A Saline solution is administered by the following:
3 to 6 liters.
2 to 3.5 hr. period.
Determine amount administered in 10 min.
Check my calculations, please.
Liter = 1.0567 quarts.
Quart = 32 oz.
3 * 1.0567 = 3.17 quarts.
3.17 * 32 = 101.44 oz.
2 hrs. = 120 min.
101.44 / 120 = .084 oz.
.084 * 10 =8.45 oz.

Answer by Alan3354(69443)

A Saline solution is administered by the following:
3 to 6 liters.
2 to 3.5 hr. period.
Determine amount administered in 10 min.
Check my calculations, please.
Liter = 1.0567 quarts.
Quart = 32 oz.
3 * 1.0567 = 3.17 quarts.
3.17 * 32 = 101.44 oz.
2 hrs. = 120 min.
101.44 / 120 = .084 oz.
.084 * 10 =8.45 oz.
==========================
Not clear. Was the period 2 hours? What was the volume?
Why did you change to ounces?